OpenStudy (adi3):

Will medal please help find the inverse of g(x) = (1)/x-1 + 3

OpenStudy (adi3):

\[\left(\begin{matrix}\\ \frac{ 1}{ x-1 }\end{matrix}\right) + 3\]

OpenStudy (adi3):

\[\frac{ 1 }{ x-1} + 3\]

OpenStudy (adi3):

if this more clear

Nnesha (nnesha):

g(x) is same as y so you can replace g(x) with y \[\huge\rm \color{ReD}{y}=\frac{ 1 }{ \color{blue}{x}-1 } +3\] to find inverse `switch x and y` \[\huge\rm \color{blue}{x}=\frac{ 1 }{ \color{red}{y}-1 } +3\] now solve for y

OpenStudy (adi3):

ok, x - 3 = \[\frac{ 1 }{ y-1 }\]

OpenStudy (adi3):

is that right

OpenStudy (adi3):

@IrishBoy123

OpenStudy (adi3):

@HELPMEPLZ!!

OpenStudy (adi3):

still on inverse, this is too much for me

Nnesha (nnesha):

sorry i was afk yes that's right \[\huge\rm x-3=\frac{ 1 }{ y-1}\] now how would you cancel out y-1 from right side ?

OpenStudy (adi3):

divde it?

imqwerty (imqwerty):

write y in terms of x and then jst switch y and x :) function which u get is inverse

Nnesha (nnesha):

its already divided by 1 (1/y-1 )to cancel it out you should do opposite of division

OpenStudy (adi3):

multiply

Nnesha (nnesha):

yes right multiply both sides by y-1

OpenStudy (adi3):

so 1x-3

Nnesha (nnesha):

multiply `both sides` by y-1 \[\huge\rm (y-1) (x-3)=\frac{ 1 }{\cancel{y-1} } \times \cancel{(y-1)}\] we can cancel out y-1 \[\large\rm (y-1)(x-3)=1\] now again to solve for y move `x-3` to the right side

OpenStudy (adi3):

ohh, then how to get y by itself?

Nnesha (nnesha):

how would you cancel x-3 from left side ?

OpenStudy (adi3):

divide

Nnesha (nnesha):

right so divide both sides by x-3 let me know what you get

Nnesha (nnesha):

it's multiplication( y-1 ) times (x-3) so yes you should divide

OpenStudy (adi3):

y-1 = 1/x-3

Nnesha (nnesha):

right whats nexxt ?

OpenStudy (adi3):

y = 1/x-3 + 1

Nnesha (nnesha):

right now we should get the common denominator